Documenta Mathematica

Band 12, 2007

Benoît Collins, James A. Mingo, Piotr Śniady, Roland Speicher Second Order Freeness and Fluctuations of Random Matrices III.

Higher order freeness and free cumulants  1–70

Marc Levine

Motivic Tubular Neighborhoods  71–146

Andreas Langer and Thomas Zink

De Rham-Witt Cohomology and Displays  147–191

Gonçalo Tabuada On the Structure of Calabi-Yau Categories

with a Cluster Tilting Subcategory  193–213

V. Gritsenko, K. Hulek and G. K. Sankaran The Hirzebruch-Mumford Volume for

the Orthogonal Group and Applications  215–241

Jan Nekovář

On the Parity of Ranks of Selmer Groups III  243–274

Jean Fasel

The Chow-Witt ring  275–312

Christian Voigt Equivariant Local Cyclic Homology

and the Equivariant Chern-Connes Character  313–359

Kiran S. Kedlaya

Erratum for “Slope Filtrations Revisited”  361–362

Richard Hill Construction of Eigenvarieties in Small Cohomological Dimensions for Semi-Simple,

Simply Connected Groups  363–397

Lionel Dorat

G-Structures Entires et Modules de Wach  399–440

Alexander Schmidt

Rings of Integers of Type K(π, 1)  441–471

Karim Johannes Becher, Detlev W. Hoffmann Isotropy of Quadratic Spaces

in Finite and Infinite Dimension  473–504

Esfandiar Nava Yazdani

Generic Observability of Dynamical Systems  505–520

Pierre Guillot

Geometric Methods for Cohomological Invariants  521–545

Igor Wigman

Erratum  547–548

Indranil Biswas and Johannes Huisman Rational Real Algebraic Models

of Topological Surfaces  549–567

Alexander Pushnitski and Grigori Rozenblum Eigenvalue Clusters of the Landau Hamiltonian

in the Exterior of a Compact Domain  569–586

William A. Stein

Visibility of Mordell-Weil Groups  587–606

Florian Ivorra Réalisation ℓ-Adique

des Motifs Triangulés Géométriques I  607–671

Dimitar P. Jetchev, William A. Stein Visibility of the Shafarevich–Tate Group

at Higher Level  673–696